OpenStudy (anonymous):

Please Help! Consider the function. f (x) = 1/2x +1 a. Find the inverse of f(x) and name it g(x). Show and explain your work. b. Use function composition to show that f(x) and g(x) are inverses of each other. c. Draw the graphs of f(x) and g(x) on the same coordinate plane. Explain how your graph shows that the functions are inverses of each other.

myininaya (myininaya):

So you are given y=x/2+1. 1st step: Solve for x. Do this and I will give you the next step.

OpenStudy (anonymous):

is it y-1=x/2?

myininaya (myininaya):

Ok you still need x on one side. How to undo that division by 2?

OpenStudy (anonymous):

to tell you the truth i really don't know. i was hoping you could help explain it

myininaya (myininaya):

To undo division, you multiply.

myininaya (myininaya):

Since you have division by 2 and you want to undo the division, then you multiply by 2. Don't forget whatever you do to one side you do to the other.